# Frequency-Guided Posterior Sampling for Diffusion-Based Image Restoration

### v1

@@ Abstract
Image restoration aims to recover high-quality images from degraded observations. When the degradation process is known, the recovery problem can be formulated as an inverse problem, and in a Bayesian context, the goal is to sample a clean reconstruction given the degraded observation. Recently, modern pretrained diffusion models have been used for image restoration by modifying their sampling procedure to account for the degradation process. However, these methods often rely on certain approximations that can lead to significant errors and compromised sample quality. In this paper, we provide the first rigorous analysis of this approximation error for linear inverse problems under distributional assumptions on the space of natural images, demonstrating cases where previous works can fail dramatically. Motivated by our theoretical insights, we propose a simple modification to existing diffusion-based restoration methods. Our approach introduces a time-varying low-pass filter in the frequency domain of the measurements, progressively incorporating higher frequencies during the restoration process. We develop an adaptive curriculum for this frequency schedule based on the underlying data distribution. Our method significantly improves performance on challenging image restoration tasks including motion deblurring and image dehazing.
@@ 1 Introduction
In the above, $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ denotes the forward operator, which we assume to be known, and $\mathbf{z}\in\mathbb{R}^{m}$ is additive Gaussian noise.
In most applications, recovering $\mathbf{x}_{0}$ is difficult because the forward operator is a many-to-one mapping. To address this challenge, researchers incorporate prior knowledge about $\mathbf{x}_{0}$, including sparsity [5, 16], low-rank structure [6, 4], total variation [31], or deep generative priors [29, 3]. One way of enforcing this knowledge is by solving a regularized optimization problem where the objective is to find an $\mathbf{x}_{0}$ that maximizes the likelihood under the chosen prior model while satisfying the constraint $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})$. Another approach is to follow the Bayesian viewpoint where instead of recovering an exact $\mathbf{x}_{0}$, we sample from $p(\mathbf{x}\mid\mathbf{y})$ where Equation 1 is the model for $p(\mathbf{y}\mid\mathbf{x}_{0})$ [36]. The advantage of this interpretation is that we can draw multiple plausible samples. In image restoration tasks, the diversity of samples can help the user pick a visually appealing reconstruction.
@@
    For linear inverse problems, for the first time in the literature, we exactly characterize the approximation gap between $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$ and commonly used approximations under distributional assumptions on the data. Further, exploiting the frequency distribution of natural images, we give precise examples of forward operators that can result in very large approximation gaps.
@@
    Our analysis guides us in developing a new unsupervised diffusion-based image restoration algorithm, denoted as Frequency Guided Posterior Sampling (FGPS). To handle the intricate relationship between the frequency characteristics of the data and the forward operator, we apply a time-varying low-pass filter to $\mathbf{y}$ in its frequency domain which progressively includes higher frequency ranges during the reverse diffusion process.
@@
    We develop a data-dependent curriculum for the low-pass filters. While being easy to implement, our algorithm leads to a significant boost in performance on challenging linear and nonlinear image restoration problems, such as motion deblurring and image dehazing.
@@ 2.1 Diffusion Models
The SDE is characterized by its drift coefficient $-\frac{1}{2}\beta(t)$ and diffusion coefficient $\sqrt{\beta(t)}$, with $\mathbf{w}_{t}$ representing the standard Wiener process over the interval $[0,1]$. The initial condition of this SDE is a datapoint $\mathbf{x}_{0}\sim p(\mathbf{x})$. The drift and diffusion coefficients are chosen such that $\mathbf{x}_{t}$ goes from the clean data distribution at $t=0$ to pure noise $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=1$. A special case is known as the Variance-Preserving SDE (VP-SDE), where we can write the transition density $p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ as
@@
where $t$ and the new Wiener process $\bar{\mathbf{w}}_{t}$ both run in reverse. A seminal result of Anderson gives that this reverse process has the same probability distribution as the forward process for every $t$ [1]. In order to estimate the time-dependent score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, diffusion models use the denoising score matching technique to train a neural network that approximates the score function [39]. Sampling from $p(\mathbf{x})$ is performed by solving a discrete-time version of Equation 4 over $T$ steps, where $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ [20, 35].
@@ 2.2 Posterior Sampling
In inverse problems, we are given measurements $\mathbf{y}$ as in Equation 1, and the goal is to recover the underlying clean $\mathbf{x}_{0}$ given $\mathbf{y}$. Because $\mathcal{A}$ is not invertible in general, the inverse problem can be highly ill-posed, and assumptions on the data are needed to make the problem tractable.
A popular Bayesian strategy to solve the problem is to develop samplers for $p(\mathbf{x}\mid\mathbf{y})$ assuming we have a prior for $p(\mathbf{x})$ modeled by a pretrained generative model such as a diffusion model. For instance we can extend Equation 4 to sample from $p(\mathbf{x}\mid\mathbf{y})$ by using a conditional score function
@@
While one could train a conditional diffusion model to approximate the conditional score, this would require retraining models for every inverse problem we wish to solve [32]. Instead, we can use Bayes rule to write the conditional score as
@@
[…]
where $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ represents the complex distribution of denoised estimates of $\mathbf{x}_{t}$. Even though $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ can be complex, the Tweedie’s estimate gives us the posterior mean $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$ as a function of the unconditional score function [17] (refer to Appendix A for more background). Thus, some unsupervised methods approximate $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ using $\bm{\mu}_{0\mid t}$, for example as a Dirac delta around $\bm{\mu}_{0\mid t}$ [9] or an isotropic Gaussian distribution with mean $\bm{\mu}_{0\mid t}$ [33]. This results in tractable approximations for the conditional score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{y})$, but introduces a potentially large approximation error. In this work, we quantify this approximation gap exactly under distributional assumptions on $\mathbf{x}_{0}$, and use our insights to develop a simple modification that reduces the approximation gap significantly.
@@ 3 Related Work
Inverse problems of the form in Equation 1 have been solved for many years with different priors, such as sparse/low-rank priors [5, 6, 4] or generative priors [3, 29]. In this work, we focus on posterior sampling methods that use pretrained diffusion models and approximate the conditional score to perform guidance [9, 33, 40, 30, 35, 8, 19]. These have been applied to inverse problems such as image deblurring, super-resolution [11], and image inpainting [27, 9]. However, they can incur significant approximation errors as we will illustrate in Section 4.1. Works such as [10, 41] propose methods to mitigate these approximation errors, but rely on strong assumptions on the data manifold. In this work, we use a more practical assumption that the data obeys a power law in the frequency domain [38] in order to develop a new approximation.
@@
A line of work has explored performing diffusion in a transformed space to perform guidance for inverse problems [23, 24]. This is in fact closely related to our work as our work is akin to performing diffusion in the Fourier domain instead, similar to [28]. Building on recent connections between diffusion models and spectral autoregression [15], our work differs from previous approaches [35, 8] by using a frequency-guided measurement schedule rather than one based on the unconditional diffusion model’s variance schedule. We refer the reader to [12] for a comprehensive survey on using diffusion models for inverse problems.
4 Analysis of Approximation Error
As illustrated in Section 2.2, the main challenge in developing unsupervised inverse problem solvers is estimating the noisy likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$, which depends on the intractable denoising distribution $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$. It is common to substitute this denoising distribution with an approximation that depends only on the posterior mean $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$, which is tractable using Tweedie’s formula. However, these methods can incur a significant approximation error by using simple distributions centered around the posterior mean instead of the true distribution.
We begin by understanding precisely when this approximation gap is high. Our analysis is inspired by the empirical observation that we can construct simple image restoration tasks where several state-of-the-art diffusion-based methods struggle to recover the underlying signal. Specifically, in Figure 1, we consider two inverse problems, both of which have forward operators that are convolutions with a given kernel. The first is when the kernel is a low-pass filter, such as a Gaussian blur kernel. The second is when the kernel is a high-pass filter, such as a Dirac kernel minus a Gaussian blur kernel. As expected, prior approaches such as Diffusion Posterior Sampling [9] and variants [10, 41] perform well at Gaussian deblurring. However, when the forward operator is a high-pass filter, remarkably these approaches all fail to produce coherent outputs. This failure appears to stem from a gap between the true conditional score and its estimate, which we demonstrate theoretically in the following section.
4.1 Theoretical Analysis
Our first contribution is a precise characterization of the approximation errors in Diffusion Posterior Sampling (DPS), a popular inverse problem solver, under specific distributional assumptions. To explain the phenomenon in Figure 1, our analysis requires two key properties.
  1. 1.
  2. 2.
For the first property, we can look at the power spectral density of the data, as defined in Equation 8, which tells us how much the signal’s power is distributed across different frequencies. For the second property, it is natural to consider a setting where the clean data $\mathbf{x}_{0}$ is drawn from a multivariate Gaussian distribution. Although these two properties seem to be at odds, the following lemma establishes an equivalence between the power spectral density of the data and its covariance.
Lemma 1.
Let $f_{k}=\frac{k}{n}$ for $k=0,\dots,n-1$ denote the DFT sample frequencies for a signal of length $n$. There exists a covariance matrix $\bm{\Sigma}_{f}$ such that for $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{f})$, the following two properties hold. First, the signal $\mathbf{x}$ follows a power law in the frequency domain with parameters $c,\beta>0$, i.e., it has a power spectral density $S(f_{k})=c|f_{k}|^{-\beta}$ for the non-zero DFT sample frequencies $f_{k}$. Second, the eigenvalues of $\bm{\Sigma}_{f}$ are precisely $c|f_{k}|^{-\beta}$.
This is a special case of a classical result known as the Wiener-Khinchin Theorem [7]. Note that this does not imply that all random vectors that follow a power law in the frequency domain are Gaussian. For example, natural image data tends to follow a radially averaged power law in the frequency domain, but is far from being distributed as a Gaussian [38]. Nonetheless, this lemma guides us in understanding the approximation gap in a setting that mimics the frequency characteristics of natural image data.
Next, we quantify the approximation gap under a multivariate Gaussian assumption on the data. Let $s^{*}(\mathbf{x}_{t},\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{y})$ be the true conditional score and let $s^{\text{DPS}}(\mathbf{x}_{t},\mathbf{y})$ be the approximation of the score utilized by DPS. Recall that the main approximation used by DPS is
$s^{\text{DPS}}(\mathbf{x}_{t},\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\bm{\mu}_{0\mid t})\,,$ (9)
where $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$. Note that $\bm{\mu}_{0\mid t}$ is a function of $\mathbf{x}_{t}$.
Theorem 2.
Suppose $\mathbf{x}_{0}$ is drawn from $\mathcal{N}(\mathbf{0},\bm{\Sigma})$ and we are given linear measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{z}$, where $\mathcal{A}(\mathbf{x}_{0})=\mathbf{A}\mathbf{x}_{0}$ and $\mathbf{z}\sim\mathcal{N}(0,\sigma_{y}^{2}I)$. Suppose that the intermediate value $\mathbf{x}_{t}$ of the continuous-time reverse diffusion from Equation 5 can be written as $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we have that
$\!\!\!\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})=(\mathbf{A}\bm{\Gamma}_{t})^{T}\!(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}\!+\sigma_{y}^{2}\bm{I})^{-1}\bm{\Delta}_{t},\!\!$ (10)
where $\bm{\Gamma}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{\Sigma}(\bar{\alpha}_{t}\bm{\Sigma}+(1-\bar{\alpha}_{t})\bm{I})^{-1}$, $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Gamma}_{t}\mathbf{x}_{t}$, $\bm{\Sigma}_{0\mid t}=\mathrm{Cov}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]=\bm{\Sigma}-\sqrt{\bar{\alpha}_{t}}\bm{\Gamma}_{t}\bm{\Sigma}$, and $\bm{\Delta}_{t}=\mathbf{y}-\mathbf{A}\bm{\mu}_{0\mid t}$. Moreover, the approximation gap $\mathcal{J}^{\text{DPS}}:=s^{\text{DPS}}(\mathbf{x}_{t},\mathbf{y})-s^{*}(\mathbf{x}_{t},\mathbf{y})$ is given by:
$\mathcal{J}^{\text{DPS}}=\left(\mathbf{A}\bm{\Gamma}_{t}\right)^{T}(\sigma_{y}^{-2}\mathbf{I}-(\mathbf{A}\bm{\Sigma}_{0\mid t}\mathbf{A}^{T}+\sigma_{y}^{2}\mathbf{I})^{-1})\bm{\Delta}_{t}.$ (11)
The proof is in Appendix B. The $\bar{\alpha}_{t}$ corresponds to the variance schedule of standard diffusion models such as DDPM [20]. Our assumption that $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ isolates the effects of the approximation gap as the generated sample is exactly $\mathbf{x}_{0}$. While we characterize this approximation gap only for DPS, this framework allows us to analytically study the gaps of many other methods, such as $\Pi$-GDM [33], since we have a formula for the noisy likelihood score. A detailed study of other works and their approximation errors would be a useful area of future work.
4.2 Empirical Study
Armed with Lemma 1 and Theorem 2, we now study the approximation gap and how the forward operator $\mathcal{A}$ interacts with $\bm{\Sigma}$ depending on how it affects certain frequency ranges. To do this, we first generate 1-D signals $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{f})$ that follow a power law in the frequency domain using the construction of $\bm{\Sigma}_{f}$ given in Lemma 1. The corresponding measurements for these signals are $\mathbf{y}=\mathbf{A}\mathbf{x}_{0}+\mathbf{z}$ with $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. As in Figure 1, we let $\mathcal{A}$ be a low-pass filter or a high-pass filter, where the low-pass filter is a Gaussian kernel of standard deviation $\sigma$, denoted as $g(\sigma)$, and the high pass filter is a Dirac kernel minus $g(\sigma)$. We then calculate the exact approximation gap between DPS and the true conditional score using Theorem 2 for different diffusion timesteps. The full experimental details of this study can be found in Appendix D.
Figure 2 shows the DPS approximation gap as a function of the diffusion timesteps (where higher timesteps correspond to a lower value of $\bar{\alpha}_{t}$) for the variance schedule used in the DDPM model for $T=1000$ [20]. We clearly see two different trends depending on the forward operator. When the operator is a low-pass filter, the DPS approximation gap is low when $t$ is either close to $T$ or close to $1$. When the operator is a high-pass filter, the DPS approximation worsens significantly as we increase $t$, likely because noise affecting higher frequency ranges in $\mathbf{x}_{t}$ is amplified by the high-pass filter. In the initial stages of the reverse process, the estimated conditional score is very inaccurate, corroborating the issues seen in Figure 1. To the best of our knowledge, this is the first time any distributional assumptions have been used to characterize the exact gap between the DPS approximation and the true conditional score. While [9] and [41] provide upper and lower bounds on this gap, known as the Jensen gap, these bounds do not exploit the structure of the data distribution and as such, can be loose.
Figure 2: We show the exact approximation gap for DPS across diffusion timesteps when the data follows a power law in the frequency domain and the forward operator is a low-pass filter or a high-pass filter of varying widths. The width denotes the $\sigma$ of the Gaussian kernel. Note the $y$-axis is in log-scale.

### v2

@@ Abstract
Image restoration aims to recover high-quality images from degraded observations. When the degradation process is known, the recovery problem can be formulated as an inverse problem, and in a Bayesian context, the goal is to sample a clean reconstruction given the degraded observation. Recently, modern pretrained diffusion models have been used for image restoration by modifying their sampling procedure to account for the degradation process. However, these methods often rely on certain approximations that can lead to significant errors and compromised sample quality. In this paper, we propose a simple modification to existing diffusion-based restoration methods that exploits the frequency structure of the reverse diffusion process. Specifically, our approach, denoted as Frequency Guided Posterior Sampling (FGPS), introduces a time-varying low-pass filter in the frequency domain of the measurements, progressively incorporating higher frequencies during the restoration process. We provide the first rigorous analysis of the approximation error of FGPS for linear inverse problems under distributional assumptions on the space of natural images, demonstrating cases where previous works can fail dramatically. On real-world data, we develop an adaptive curriculum for our method’s frequency schedule based on the underlying data distribution. FGPS significantly improves performance on challenging image restoration tasks including motion deblurring and image dehazing. The source code is available at https://github.com/darshanthaker/FGPS.
@@ 1 Introduction
In the above, $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ denotes the forward operator, which we assume to be known, and $\mathbf{z}\in\mathbb{R}^{m}$ is additive Gaussian noise. In most applications, recovering $\mathbf{x}_{0}$ is difficult because the forward operator is a many-to-one mapping. To address this challenge, researchers incorporate prior knowledge about $\mathbf{x}_{0}$, including sparsity [5, 16], low-rank structure [6, 4], total variation [32], or deep generative priors [29, 3]. One way of enforcing this knowledge is by solving a regularized optimization problem where the objective is to find an $\mathbf{x}_{0}$ that maximizes the likelihood under the chosen prior model while satisfying the constraint $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})$. Another approach is to follow the Bayesian viewpoint where instead of recovering an exact $\mathbf{x}_{0}$, we sample from $p(\mathbf{x}\mid\mathbf{y})$ where Equation˜1 is the model for $p(\mathbf{y}\mid\mathbf{x}_{0})$ [38]. The advantage of this interpretation is that we can draw multiple plausible samples. In image restoration tasks, the diversity of samples can help the user pick a visually appealing reconstruction.
@@
    We develop a new unsupervised diffusion-based image restoration algorithm, denoted as Frequency Guided Posterior Sampling (FGPS). Our algorithm leverages the observation that diffusion models generate images hierarchically in the frequency domain [30]. Thus, to control the approximation error from estimating $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}\mid\mathbf{x}_{t})$, we apply a time-varying low-pass filter to $\mathbf{y}$ which progressively includes higher frequency ranges during the reverse diffusion process.
@@
    We then perform a precise theoretical analysis of the approximation errors in posterior sampling methods for linear inverse problems under distributional assumptions on the data. Next, exploiting the frequency distribution of natural images, we give precise examples of forward operators where FGPS has small approximation error compared to prior work.
@@
    In practice, we develop a data-dependent curriculum for the low-pass filters which allows the reverse process to adapt to spectral properties of the data distribution. While being easy to implement, our algorithm leads to a significant boost in performance on challenging linear and nonlinear image restoration problems, such as motion deblurring and image dehazing.
@@ 2.1 Diffusion Models
The SDE is characterized by its drift coefficient $\frac{1}{2}\beta(t)$ and diffusion coefficient $\sqrt{\beta(t)}$, with $\mathbf{w}_{t}$ representing the standard Wiener process over the interval $[0,1]$. The initial condition of this SDE is a datapoint $\mathbf{x}_{0}\sim p(\mathbf{x})$. The drift and diffusion coefficients are chosen such that $\mathbf{x}_{t}$ goes from the clean data distribution at $t=0$ to pure noise $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ at $t=1$. A special case is known as the Variance-Preserving SDE (VP-SDE), where we can write the transition density $p_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ as
@@
where $t$ and the new Wiener process $\bar{\mathbf{w}}_{t}$ both run in reverse. A seminal result of Anderson gives that this reverse process has the same probability distribution as the forward process for every $t$ [1]. In order to estimate the time-dependent score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, diffusion models use the denoising score matching technique to train a neural network $s_{\theta}(\cdot)$ that predicts the noise added to a datapoint, which approximates the score function [40]. Sampling from $p(\mathbf{x})$ is performed by solving a discrete-time version of Equation˜4 over $T$ steps, where $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ [20, 37].
@@ 2.2 Posterior Sampling
In inverse problems, we are given measurements $\mathbf{y}$ as in Equation˜1, and the goal is to recover the underlying clean $\mathbf{x}_{0}$ given $\mathbf{y}$. Throughout the paper, we will refer to $\mathcal{A}$ as the known forward operator. Because $\mathcal{A}$ is not invertible in general, the inverse problem can be highly ill-posed, necessitating data modelling assumptions.
A popular Bayesian strategy to solve the problem is to develop samplers for $p(\mathbf{x}\mid\mathbf{y})$ assuming we have a prior for $p(\mathbf{x})$ modeled by a pretrained generative model such as a diffusion model. For instance we can extend Equation˜4 to sample from $p(\mathbf{x}\mid\mathbf{y})$ by using a conditional score function
@@
While one could train a conditional diffusion model to approximate the conditional score, this would require retraining models for every inverse problem we wish to solve [33]. Instead, we use Bayes rule to write the conditional score as
@@
$p(\mathbf{y}\mid\mathbf{x}_{t})=\int p(\mathbf{y}\mid\mathbf{x}_{0})p(\mathbf{x}_{0}\mid\mathbf{x}_{t})d\mathbf{x}_{0},$ (7)
where $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ represents the complex distribution of denoised estimates of $\mathbf{x}_{t}$. Even though $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ can be complex, the Tweedie’s estimate gives us the posterior mean $\bm{\mu}_{0\mid t}=\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t}]$ as a function of the unconditional score function [17] (refer to Appendix A for more background). Thus, some unsupervised methods approximate $p(\mathbf{x}_{0}\mid\mathbf{x}_{t})$ using $\bm{\mu}_{0\mid t}$, for example as a Dirac delta around $\bm{\mu}_{0\mid t}$ [9] or an isotropic Gaussian distribution with mean $\bm{\mu}_{0\mid t}$ [35]. This results in tractable approximations for the conditional score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{y})$, but introduces a potentially large approximation error. In this work, we propose a novel approximation for the conditional score and quantify its approximation gap exactly under distributional assumptions on $\mathbf{x}_{0}$.
@@ 3 Related Work
Inverse problems of the form in Equation˜1 have been solved for many years with different priors, such as sparse/low-rank priors [5, 6, 4] or generative priors [3, 29]. In this work, we focus on posterior sampling methods that use pretrained diffusion models and approximate the conditional score to perform guidance [9, 35, 41, 31, 37, 8, 19]. These have been applied to inverse problems such as image deblurring, super-resolution [11], and image inpainting [27, 9]. However, they can incur significant approximation errors as we will illustrate in Section˜5. Works such as [10, 42] propose methods to mitigate these approximation errors, but rely on strong assumptions on the data manifold. In this work, we use a more practical assumption that the data obeys a power law in the frequency domain [34] in order to develop a new approximation.
@@
A line of work has studied diffusion models in a transformed space to perform guidance for inverse problems [23, 24]. This is in fact closely related to our work as our work is akin to performing diffusion in the Fourier domain instead, similar to [28]. Building on recent connections between diffusion models and spectral autoregression [15], our work differs from previous approaches [37, 8] by using a frequency-guided measurement schedule rather than one based on the unconditional diffusion model’s variance schedule. We refer the reader to [12] for a comprehensive survey on using diffusion models for inverse problems.